Smith borrows $440 for five months at 11% simple interest. How much interest is paid?
Answer by mananth(16946)

I=p*r*t
p=440
r=11%
t=5/12 years
...
I=440*0.11*5/12
I = $ 20.16
...
m.ananth@hotmail.ca